### v1

# ProactiveMobile: A Comprehensive Benchmark for Boosting Proactive Intelligence on Mobile Devices

Abstract
Multimodal large language models (MLLMs) have made significant progress in mobile agent development, yet their capabilities are predominantly confined to a reactive paradigm, where they merely execute explicit user commands. The emerging paradigm of proactive intelligence, where agents autonomously anticipate needs and initiate actions, represents the next frontier for mobile agents. However, its development is critically bottlenecked by the lack of benchmarks that can address real-world complexity and enable objective, executable evaluation. To overcome these challenges, we introduce ProactiveMobile, a comprehensive benchmark designed to systematically advance research in this domain. ProactiveMobile formalizes the proactive task as inferring latent user intent across four dimensions of on-device contextual signals and generating an executable function sequence from a comprehensive function pool of 63 APIs. The benchmark features over 3,660 instances of 14 scenarios that embrace real-world complexity through multi-answer annotations. To ensure quality, a team of 30 experts conducts a final audit of the benchmark, verifying factual accuracy, logical consistency, and action feasibility, and correcting any non-compliant entries. Extensive experiments demonstrate that our fine-tuned Qwen2.5-VL-7B-Instruct achieves a success rate of 19.15%, outperforming o1 (15.71%) and GPT-5 (7.39%). This result indicates that proactivity is a critical competency widely lacking in current MLLMs, yet it is learnable, emphasizing the importance of the proposed benchmark for proactivity evaluation.
1 Introduction
Figure 1: A comparison of proactive and reactive paradigm in mobile agents.
Fueled by rapid advancements in MLLMs [47, 49, 3], mobile agents have achieved substantial breakthroughs[46, 14] such as interface comprehension [12, 56], conversational interaction [36, 21], and task planning [8, 37].
However, these agents share a fundamental constraint: they are confined to a reactive paradigm, functioning as passive executors of direct user commands [34, 10]. These models places the entire cognitive burden on the user, from need identification to goal articulation [30], thereby relegating the agent to the role of a high-level tool and fundamentally limiting its potential for seamless integration into daily life [23].
The limitations of the reactive paradigm are becoming a critical bottleneck, propelling a fundamental shift towards proactive intelligence—the undisputed next frontier for mobile agents. This vision represents not merely an incremental improvement, but a complete re-imagining of the agent’s role: rather than being a passive tool, it evolves into a genuinely helpful assistant by autonomously anticipating user needs and initiating actions [26, 43, 44]. The profound implication is a future of human-agent collaboration where cognitive burden is minimized, and interaction feels seamlessly intuitive [4, 43, 30]. Recognizing this transformative potential, pioneering studies have indeed validated the core premise of proactivity[9, 26, 43, 44].
Despite these promising initial steps, the current research landscape for proactive agents remains fragmented and lacks a unified foundation. A core deficiency is that existing benchmarks [26, 44] oversimplify the task: they rely on abstracted contexts and crucially assume a single “correct” action per scenario. This ignores the inherent subjectivity and diversity of user preferences, forcing the complex one-to-many mapping of proactive suggestions into an unrealistic one-to-one paradigm. This flawed premise is exacerbated by the metrics used for evaluation. For instance, ProactiveAgent’s [26] binary reward model is too coarse to differentiate partial from complete failures, while FingerTip-20K [44] relies on cosine similarity, which captures semantic relevance but ignores functional correctness and executability. Beyond definition and evaluation, a third major shortcoming lies in the output format. Both benchmarks rely on generating natural language recommendations, a format that is inherently ambiguous and lacks a direct path to on-device execution, creating a critical gap between suggesting a task and actually performing it. This confluence of issues (an ill-defined task rooted in oversimplification, superficial evaluation, and a non-executable output format) critically bottlenecks the systematic advancement of the field.
To address these critical gaps, we introduce ProactiveMobile, a comprehensive benchmark designed to systematically advance research on proactive mobile agents. To mitigate oversimplification, ProactiveMobile formalizes the proactive task by requiring agents to predict actions based on four dimensions of on-device contextual signals: user profile, device status, world information, and behavioral trajectories. Figure 1 clearly illustrates this process and contrasts it with the reactive agent paradigm. To tackle the ignorance of user preferences, ProactiveMobile embraces the one-to-many nature of proactivity: each instance is annotated and manually verified with one to three target actions. The resulting benchmark is substantial, comprising 3,660 instances of 14 distinct scenes spanning a diverse range of real-world scenarios. Furthermore, to overcome the inherent ambiguity and subjectivity of evaluating natural language suggestions, we introduce a crucial constraint: models must translate their intents into executable actions. We achieve this by constructing a comprehensive function pool of 63 APIs, requiring models to output specific function sequences. This approach transforms the evaluation from a subjective text-matching problem into an objective, structured task.
To establish baselines on our benchmark, we fine-tuned Qwen2.5-VL-7B-Instruct [3] and MiMo-VL-7B-SFT-2508 [35] on the training set. We then evaluated its performance on ProactiveMobile alongside a suite of leading closed-source models, including o1 [17], GPT-5 [29], and Gemini-2.5-Pro [7]. Our fine-tuned Qwen2.5-VL-7B-Instruct achieves a success rate of 19.15% on the exact function sequence matching, significantly outperforming closed‑source models, including o1 (15.71%), GPT-5 (7.39%), and Gemini-2.5-Pro (8.91%).
These results offer two critical insights. First, this superior performance supports our hypothesis: proactivity is a specialized capability that requires targeted, domain-specific training, as provided by ProactiveMobile. Even the most powerful general-purpose models fail to master it out-of-the-box. Second, although proactivity is learnable, the performance of trained models still fails to meet the requirements for on-device deployment. This indicates that proactive intelligence is a highly challenging research problem, which in turn underscores the significance of our work and the necessity of the proposed benchmark. Our contributions are as follows:
  1. 1.
    We propose a novel and comprehensive task formalization for proactive mobile agents, grounding the problem in rich, multi-dimensional real-world context.
  2. 2.
    We construct and open-source ProactiveMobile, comprising 3,660 multi-intent instances across 14 scenarios. To facilitate deployment and fine-grained evaluation, all intents are mapped to corresponding function sequences via a predefined function pool of 63 APIs.
  3. 3.
    We provide an in-depth empirical analysis, establishing strong baselines and revealing that proactivity is a specialized capability lacking in current general models, thereby highlighting critical challenges and future research directions. Notably, we will release our model weights to foster progress within the research community.
2 Related Work
2.1 LLM-Based Mobile Interaction
The advent of MLLMs has ushered in a new era of mobile agents capable of understanding natural language instructions and visual UI elements to perform actions autonomously. These LLM-based mobile agents represent a paradigm shift, enabling users to accomplish intricate, multi-step tasks on web, mobile, or desktop applications via simple conversational commands [48, 34].
A core capability of these agents is GUI understanding, or GUI grounding, where MLLMs interpret screen layouts by combining visual perception with textual information. To enhance this, specialized models [41, 45, 38] and methods [12, 56, 33, 6] have been developed to better process GUI-specific modalities. The rapid progress in this area is also fueled by the development of specialized datasets, including large-scale annotated datasets [41, 15, 20, 19, 24] and data pipelines [45, 19, 24].
Another critical area is task planning and execution. LLMs excel at decomposing high-level natural language commands into a series of executable actions. However, methods based on static prompting often struggled with long-horizon tasks and dynamic environments [51, 34, 42, 40]. Some researches explore fine-tuning or reinforcement learning to enhance the reasoning and prediction capabilities of MLLMs in relative tasks [39, 53, 28, 27, 13]. The maturation of the field is also marked by comprehensive benchmarks [25, 50, 54, 55, 53], which provide standardized environments for evaluating agent performance on realistic tasks.
2.2 Proactive Agents
The paradigm of intelligent agents is undergoing a significant shift, moving from reactive systems that await explicit user commands to proactive agents that anticipate user needs [9]. By inferring likely intentions and preemptively offering or executing useful actions, these agents can enhance user engagement and task efficiency [32]. Research in proactivity has evolved through several distinct stages.
Initial explorations in this domain have largely focused on proactive conversational agents. Instead of passively responding, these systems actively guide the dialogue by asking clarifying questions, suggesting relevant topics, or steering the conversation towards a productive goal [22, 9, 10]. While foundational, their proactivity is primarily confined to the conversational level.
Building on this, subsequent research has delved deeper into proactive intent inference, where the agent’s goal is to predict a user’s next action or ultimate goal from their behavior. These approaches can be broadly categorized into two types: those that explicitly prompt the user for clarification to confirm intent [31, 52], and those that implicitly infer intent from contextual cues and behavioral history [18, 44]. This line of work is crucial for understanding user needs before they are articulated.
The most advanced form of proactivity involves agents that not only anticipate needs but also autonomously execute or propose complete tasks. This represents the ultimate goal of delivering value to the user with minimal friction. However, existing work in this advanced stage often faces significant limitations. Some studies are confined to narrow, specific domains like smart home control, limiting their generalizability [5]. Others predict overly simplistic, single-step tasks, often within simulated or artificial scenarios that do not capture the nuances of genuine user interactions [26, 44]. Our work addresses these gaps by introducing a benchmark where the data is deeply grounded in diverse, realistic scenarios, designed to evaluate an agent’s ability to recommend complex, multi-step tasks.
3 Benchmark
Figure 2: The overview of data generation.
In this section, we define the task and detail the benchmark construction process. Due to space limitations, we provide comprehensive implementation details in the Appendix, including the prompt templates used for data generation, the design of the annotation platform, annotator training materials, annotation guidelines, quality control procedures, and other technical specifications.
3.1 Task Definition
We define proactive intelligence as the task of predicting user’s latent intentions based on their user profile, device status, world information, and behavioral trajectories. The details of these four categories are as follows:
  • User Profile. The user’s static attributes and dynamic behavioral characteristics encompass basic information, long-term behavioral habits, and personal preferences.
  • Device Status. Real-time device and immediate environmental states include hardware, battery level, network status, location, and notifications.
  • World Information. External circumstances, including weather, time of day, and public holidays.
  • Behavioral Trajectories. A temporal sequence of user-device interactions that reveals evolving intent.
In terms of representation, user profile, device status, and world information are expressed in natural language, while behavioral trajectories are represented either as textual descriptions or sequences of GUI screenshots. To facilitate command execution, all intents are mapped into a unified sequence of executable functions. Consequently, the complete proactive intelligence task can be formalized as:
$\mathcal{T}=\{(\boldsymbol{I}_{k},\boldsymbol{F}_{k})\}_{k=1}^{a}=\text{Predict}(\mathbf{U},\mathbf{D},\mathbf{W},\mathbf{B}).$ (1)
Here, $\mathbf{U}$, $\mathbf{D}$, $\mathbf{W}$, and $\mathbf{B}$ represent the user profile, device status, world information, and behavioral trajectories at the decision moment. For each decision point, there may exist multiple valid intent–function pairs, denoted as the ground-truth set $\mathcal{T}$. The model generates a single predicted pair:
$(\hat{\boldsymbol{I}},\hat{\boldsymbol{F}})=\boldsymbol{M}_{\theta}(\mathbf{U},\mathbf{D},\mathbf{W},\mathbf{B}),$ (2)
[…]
where $\hat{\boldsymbol{F}}$ is non-empty only if $\hat{\boldsymbol{I}}$ is actionable and can be mapped to at least one function from the predefined function pool $\mathbb{F}$, i.e., $\hat{\boldsymbol{F}}\subseteq\mathbb{F}$; otherwise, $\hat{\boldsymbol{F}}$$\varnothing$.
The prediction is considered correct if the model output matches any ground-truth pair:
$(\hat{\boldsymbol{I}},\hat{\boldsymbol{F}})\in\mathcal{T}.$ (3)
3.2 Dataset Construction
This section is organized into three main sections. First, we elaborate on the acquisition methods for behavioral trajectories. Second, we outline the end-to-end data generation process. Finally, we provide a detailed account of the data auditing mechanism.
3.2.1 Acquisition of Behavioral Trajectories
User behavioral trajectories serve as the foundation for intent prediction. In cases where direct access to user actions is unavailable, screenshots are used as substitutes. We define these two modalities as:
  • Multimodal trajectories: Sequences of mobile screenshots captured during user interactions, combined with corresponding text commands from both public and self-built datasets, summarized in Table 1.
  • Text trajectories: Textual logs of user actions are derived from GUI traces. We first categorize and deduplicate text commands, then employ Claude-Sonnet-4 to generate text-based action trajectories via prompt-based expansion.
Dataset Description Usage
GUI-Odyssey [25] A dataset for cross-App mobile GUI navigation 21,669
AITZ [50] Largest dataset in android GUI navigation field 9,413
CAGUI [54] A real-world Chinese android GUI benchmark 3,196
MobileAgentBench Self-collected GUI data from mobile devices 12,481
Table 1: Summary of GUI dataset.
3.2.2 Generation Pipeline
The data generation process involves five key steps, as illustrated in Figure 2.
  1. 1.
    Generate contextual information. Based on user behavioral trajectories and relevant commands, we randomly employ Claude-Sonnet-4[2], Gemini-2.5-Pro [7], and GPT-5 [29] to generate three complementary information components—user profile, device status, and world information—thereby constructing a comprehensive contextual information stream. The generated information then undergoes a plausibility check using the o1 [17]. If the content is deemed implausible, it is discarded and subsequently regenerated.
  2. 2.
    Generate potential intentions. Leveraging the contextual information, multiple MLLM models simulate users’ potential next-step intentions. These intentions represent tasks that agents can recommend proactively, triggered by specific conditions and personalized according to the user profile. To ensure both diversity and quality of generated intentions, we select six state-of-the-art closed-source MLLMs (Claude-Sonnet-4 [2], Claude-Sonnet-3.7 [1], Gemini-2.5-Pro [7], Gemini-2.0-Flash [11], o1 [17], and GPT-4o [16]) with strong multimodal understanding and reasoning capabilities. To unify their outputs, Gemini-2.5-Flash [7] is prompted to semantically cluster the 30 generated candidates and extract the top-3 representative intentions (cluster centroids) ranked by overall support across models.
  3. 3.
    Add interfering information. To enhance model robustness, we intentionally inject irrelevant textual noise into the user profile, device states, and environmental information. The injected noise consists of task-irrelevant yet semantically coherent text generated by Gemini-2.5-Pro [7] through carefully designed prompts. This process preserves overall logical consistency while training the model to focus on salient and task-relevant signals. On average, the amount of injected noise is approximately 5–20 times the volume of task-relevant information.
  4. 4.
    Map to function. Convert intentions into function calls. We uniformly convert textual instructions generated by multiple MLLMs into executable function-call sequences. This conversion is performed by Claude-Sonnet-4 [2], which is prompted to select appropriate functions from a predefined function pool to fulfill each recommended task. The resulting sequence may include one or more functions, while a zero-function sequence indicates that no action is required and triggers the no-recommendation logic.
  5. 5.
    Three-stage review. A three-stage review mechanism—comprising rule-based checks, agent evaluations, and expert reviews—is adopted to filter and validate generated data, ensuring reliability and accuracy.
3.2.3 Three-Stage Review
To ensure data quality, we implement a comprehensive quality control process spanning three stages: rule-based filtering, agent evaluation, and expert review.
  1. 1.
    Rule-based filtering. Automatically removes entries that fail to meet format and consistency requirements.
  2. 2.
    Agent evaluation. We employ Gemini-2.5-Pro [7] to assess the internal consistency among textual information, action trajectories, and recommended actions. Using a prompt-based evaluation framework, the model examines textual information for authenticity and naturalness, trajectories for realism and temporal coherence, and recommendations for contextual appropriateness and executability.
  3. 3.
    Expert review. Experts verify the remaining entries for factual accuracy, internal logic feasibility, and action feasibility. A team of 30 trained annotators, each with prior experience in human–computer interaction and data annotation, conducts the verification process. All annotators undergo standardized training and trial labeling sessions to align annotation criteria and resolve ambiguities. To ensure data quality, each data point is independently annotated by three annotators. An item is considered valid and accepted for the final dataset only if at least two annotators are in agreement on its label. This extensive cleaning and correction process represents a four-month effort with a total investment of \$210,000. Throughout this period, experts collaboratively refine and validate the dataset to ensure its reliability and consistency.
3.3 Function Pool Construction
To facilitate on-device execution and standardized evaluation, we transformed textual instructions into a unified function call format by creating a predefined function pool. Our construction process involved a multi-stage pipeline. First, we manually categorized instructions into 14 distinct scenes. Then, we employed LLMs to initially generate function sequences and subsequently refine them by merging similar functions and parameters while pruning infrequent ones. Following this automated phase, we defined a formal schema for each function, annotating parameter data types and specifying required arguments. Finally, the entire function pool underwent a rigorous manual verification by five experienced doctoral researchers specializing in AI agents and system design, who cross-checked all definitions to ensure semantic consistency, correctness, and overall coherence.
3.4 Difficulty Definition
To systematically evaluate model performance across different levels of challenge, we establish a three-tier difficulty system. We classify each data item based on the number of correct predictions from a panel of five powerful models: Claude-Sonnet-4 [2], Claude-Sonnet-3.7 [1], GPT-4o [16], Gemini-2.5-Pro [7], and Gemini-2.5-Flash [7]. The difficulty level is defined as follows:
  • Level 1: 4–5 models correct.
  • Level 2: 2–3 models correct.
  • Level 3: 0–1 models correct.
To validate this automatic classification, a group of five experienced doctoral researchers independently assessed a stratified sample of data items. The resulting difficulty annotations showed an inter-rater agreement of over 95% with our model-based difficulty levels, confirming the reliability and consistency of the proposed three-tier system.
Finally, in Figure 3 and Table 2, we illustrate the dataset information for our benchmark.
Figure 3: Distribution of the 14 primary user intent categories, demonstrating the benchmark’s broad scenario coverage (e.g., Personal Management, Office Work, Food & Dining).
Split Data Type Scenes Items Intents Images Ave. Image Functions Ave. Functions L1 L2 L3
Train Multimodal 12 4,438 8,977 32,418 7.30 9,964 1.11 308 1,376 2,754
Text 12 4,438 4,438 - - 8,259 1.86 372 1,208 2,858
Test Multimodal 14 1,832 3,711 14,341 7.83 4,173 1.24 118 613 1,101
Text 14 1,828 2,676 - - 2,266 0.85 259 711 858
Table 2: Statistics of the ProactiveMobile dataset, broken down by Train and Test splits and data modality. The table details the composition of our benchmark, including the number of scenes, items, intents, functions, and distribution of different difficulties. Notably, the test set includes two additional scenes (14 vs. 12) not present in the training set, which form our dedicated out-of-distribution (OOD) evaluation split.
4 Experiments
This section highlights the value and contributions of our work by benchmarking against state-of-the-art closed-source MLLMs.
Model     L1     L2
    Multimodal Text All     Multimodal Text All
    SR FTR SR FTR SR FTR     SR FTR SR FTR SR FTR
GPT-5     5.08 69.23 20.46 15.07 15.65 20.82     6.20 61.47 16.34 27.29 11.64 33.74
GPT-4o     11.02 95.24 18.53 42.33 16.18 47.03     5.87 92.31 9.14 54.49 7.63 60.77
o1     18.64 46.88 27.41 2.71 24.67 8.30     11.58 32.58 21.94 4.64 17.14 10.32
Gemini-2.5-Pro     6.78 66.67 10.81 62.87 9.55 63.32     11.91 67.44 7.03 80.91 9.29 78.38
Qwen2.5-VL-7B     0.85 80.00 2.32 50.00 1.86 54.84     1.14 71.05 2.39 65.81 1.81 67.10
MiMo-VL-7B-SFT     5.08 42.10 7.34 53.49 6.63 51.43     4.08 62.12 5.63 53.88 4.91 55.79
Qwen2.5-VL-7B+Proactive     19.49 38.23 37.07 6.85 31.56 11.07     14.52 24.31 29.39 10.08 22.51 13.18
MiMo-VL-7B-SFT+Proactive     15.25 27.27 20.08 51.02 18.57 47.60     11.42 24.03 16.60 50.12 14.20 44.00
Model     L3     Avg
    Multimodal Text All     Multimodal Text All
    SR FTR SR FTR SR FTR     SR FTR SR FTR SR FTR
GPT-5     5.09 44.80 8.28 37.55 6.48 41.04     5.46 51.17 8.67 27.38 7.39 34.20
GPT-4o     4.27 90.45 4.66 52.19 4.44 70.06     5.24 91.27 8.37 51.03 6.80 61.67
o1     11.90 19.75 14.45 10.06 13.02 14.76     12.23 25.05 19.20 5.95 15.71 11.87
Gemini-2.5-Pro     9.26 66.82 7.58 82.54 8.53 74.03     9.99 66.96 7.82 76.54 8.91 73.61
Qwen2.5-VL-7B     0.73 71.05 1.28 58.06 0.97 65.22     0.87 71.77 1.86 60.17 1.37 64.22
MiMo-VL-7B-SFT     2.73 58.02 2.80 42.48 2.76 50.82     3.33 57.87 4.54 50.72 3.93 53.10
Qwen2.5-VL-7B+Proactive     13.17 24.14 16.20 11.85 14.50 17.74     14.03 25.15 24.29 9.99 19.15 14.77
MiMo-VL-7B-SFT+Proactive     11.99 27.12 12.24 35.37 12.10 30.87     12.01 26.26 15.04 46.12 13.53 39.24
Table 3: Overall performance comparison of our fine-tuned model (+Proactive) against baselines on the ProactiveMobile test set. We report two key metrics: Success Rate (SR), where higher is better, and False Trigger Rate (FTR), where lower is better. The comparison is broken down by task difficulty (L1-L3) and data modality. For each metric, the best result is in bold and the second-best is underlined. All scores are in percentage (%).
4.1 Setting
Fine-tuned Model. To create a specialized proactive agent, we perform full-parameter supervised fine-tuning (SFT) on Qwen2.5-VL-7B-Instruct [3] and MiMo-VL-7B-SFT-2508 [35]. A core aspect of our methodology is the defined output format: the model was trained to co-generate both a natural language recommendation instruction and the corresponding executable function sequence. We utilize the 8,876 instances from the training split of ProactiveMobile. Further details regarding data pre-processing, specific hyperparameters, and the hardware environment are provided in Appendix.
Baseline Models. To benchmark against the current state-of-the-art, we evaluate several leading proprietary MLLMs, including GPT-5 [29], GPT-4o [16], Gemini-2.5-Pro [7], o1 [17], unfinetuned Qwen2.5-VL-7B-Instruct [3], and unfinetuned MiMo-VL-7B-SFT-2508 [35]. All baseline models were evaluated in a zero-shot setting. To ensure a fair comparison, the standardized prompt instructed these models to adopt the same output format. We design a standardized prompt that provided each model with the same multi-dimensional context (user profile, device status, etc.) and the list of available functions from our API pool, instructing them to output the appropriate function call sequence.
4.2 Metrics
Evaluating proactive intelligence presents unique challenges, especially given the one-to-many nature of valid actions in ProactiveMobile, where a single context can map to multiple ground-truth sequences. A naive evaluation metric would either be too brittle (penalizing functionally correct but formally different predictions) or too lenient. To address this, we define two core metrics, Success Rate and False Trigger Rate, whose final values are determined by a sophisticated evaluation protocol designed specifically for this one-to-many context, as detailed below.
Training Strategy Multimodal Text All
SR FTR SR FTR SR FTR
Func. 11.57 100.00 5.31 100.00 8.44 100.00
Rec.+Func.(ours) 14.03 25.15 24.29 9.99 19.15 14.77
Think+Func. 7.92 100.00 3.77 99.83 5.85 99.89
Think+Rec.+Func. 4.97 3.02 9.79 1.80 7.38 2.21
Table 4: Ablation study on the impact of different output formats. We compare our primary Text Recommendation + Function strategy against variants that only output the Function, or include an additional reasoning (Think) step.
Model SR FTR
GPT-5 10.29 46.15
GPT-4o 3.13 52.63
o1 18.75 3.57
Gemini-2.5-pro 12.50 54.17
Qwen2.5-VL-7B-Instruct 1.56 66.67
MiMo-VL-7B-SFT-2508 4.69 45.45
Qwen2.5-VL-7B-Instruct + Proactive 15.63 27.59
MiMo-VL-7B-SFT-2508 + Proactive 7.81 40.91
Table 5: Performance on the Out-of-Distribution (OOD) test set. This set comprises 64 instances from two scenarios (Logistics Delivery and Smart Home) that were entirely absent from the training data.
4.2.1 Core Metrics
1. Success Rate (SR). This is our primary, binary success metric, designed to measure perfect functional equivalence. A prediction is considered accurate not based on simple string comparison, but on whether it is semantically and functionally identical to a valid ground truth. To make this judgment, we employ a powerful LLM judge (Gemini-2.5-pro [7])  [^1] To ensure the validity of judgment, we verified the consistency between the model and human experts, and achieved a consistency of 98%.. An instance receives a final SR score of 1 only if the model’s prediction is deemed functionally equivalent to one of the valid ground-truth answers; otherwise, it is 0. Given ProactiveMobile’s one-to-many nature, the precise procedure for selecting the “best” ground truth to compare against is critical, and is elaborated in our Best-Match Selection Protocol (Section 4.2.2).
2. False Trigger Rate (FTR). This metric measures the model’s reliability in non-trigger scenarios. It quantifies the rate at which the model incorrectly generates an action when the ground truth specifies that no action should be taken. Let $N_{no-action}$ be the total number of instances where the ground-truth set is empty ($G={\emptyset}$), and $N_{ft}$ be the number of those instances where the model falsely triggers a non-empty $S_{pred}$. The FTR is calculated as: $\text{FTR}=\frac{N_{ft}}{N_{no-action}}$
4.2.2 Best-Match Selection Protocol
The aforementioned protocol dictates how a model’s prediction ($S_{pred}$) is scored against the set of ground-truth candidates ($G={S_{label,1},...}$) to yield the final SR score. It is a two-stage process designed to be both rigorous and fair:
Stage 1: Prioritize Perfect Functional Equivalence. We first check if the model’s prediction is functionally equivalent (as judged by our LLM referee) to any of the ground-truth sequences. If one or more such ”perfect matches” are found, the SR for this instance is immediately set to 1, and the protocol terminates for this instance. One of these perfect matches is randomly selected as the best match ($S_{label}^{*}$) for any further analysis.
Stage 2: F1-Score Fallback for Imperfect Predictions. If no perfect match is found in Stage 1, the SR for this instance is definitively 0. However, for consistent and fair analysis, we still need to select a single “closest” ground truth. In this scenario, we identify the ground-truth candidate that maximizes the F1-score (calculated on the sets of function names) when compared with $S_{pred}$. To compute this score, we treat both the prediction and the ground truth as unordered sets of function names, thus ignoring parameters and sequence order. This allows us to calculate the harmonic mean of precision (the fraction of predicted functions that are correct) and recall (the fraction of correct functions that were predicted). This F1-maximizing sequence is then designated as the best match ($S_{label}^{*}$).
Why this protocol? This two-stage design serves a crucial purpose. It establishes perfect functional correctness as the unambiguous gold standard for success, which is directly reflected in our primary SR metric. The F1-fallback mechanism, meanwhile, ensures a robust and consistent process for handling failures, providing a fair basis for comparison and deeper analysis even when the primary success condition is not met.
4.3 Overall Performance
Table 3 presents a comprehensive performance analysis, revealing several critical insights into the current landscape of proactive intelligence.
Fine-tuning on ProactiveMobile consistently unlocks SOTA capabilities. The most striking result is the significant impact of fine-tuning on our benchmark. This effect is consistent across different base models: fine-tuning boosts the Qwen2.5-VL-7B-Instruct from a mere 1.37% to a state-of-the-art 19.15% Success Rate, and similarly elevates the MiMo-VL-7B-SFT-2508 from 3.93% to 13.53%. Our fine-tuned Qwen model establishes a new benchmark, significantly outperforming the top-performing proprietary model, o1 (15.71% SR). This substantial gap unequivocally demonstrates that proactivity is a specialized, learnable skill requiring domain-specific adaptation, validating ProactiveMobile as an essential training resource.
Multimodal reasoning remains a key bottleneck. The performance disparity across data types reveals a core challenge. For our top-performing model (Qwen2.5-VL-7B + Proactive), the SR on Text tasks (24.29%) is substantially higher than on Multimodal tasks (14.03%). This performance delta suggests that grounding abstract intents within noisy, real-world GUI screenshots introduces significant complexity, highlighting robust visual comprehension as a critical area for future advancement in on-device proactive intelligence.
The low absolute scores validate the task’s inherent difficulty. Despite the strong relative performance of our fine-tuned model, the absolute SR scores remain modest across the board. The fact that the state-of-the-art sits just under 20% confirms that reliable, functionally correct proactive intelligence is a profoundly difficult and unsolved problem. This finding validates ProactiveMobile not as a benchmark for a saturated task, but as a challenging and indispensable testbed designed to catalyze genuine breakthroughs in the field.
4.4 Generalization to Out-of-Distribution Scenarios
To assess generalization, we evaluated all models on an out-of-distribution (OOD) test set comprising 64 instances from two scenarios—Logistics Delivery and Smart Home—that were entirely absent from the training data. The results in Table 5 reveal a telling dichotomy. On one hand, o1 emerges as the top performer with an 18.75% SR, likely leveraging its vast pre-training to handle novel concepts. On the other hand, our fine-tuned Qwen2.5-VL-7B-Instruct + Proactive model secures a strong second place at 15.63% SR, significantly outperforming other powerful generalists like Gemini-2.5-Pro (12.50%), GPT-5 (10.29%), and GPT-4o (3.13%). This demonstrates that while immense scale offers one path to generalization, our fine-tuning approach effectively imparts a more robust and transferable understanding of proactive logic. It validates that the skills learned on ProactiveMobile are not mere pattern matching, but represent a promising step toward truly generalizable proactive intelligence.
4.5 Ablation Study
To validate our training and output format (Recommendation + Function), we conducted an ablation study comparing it with variants that either omitted the recommendation or added an explicit Think step. The results in Table 4 are decisive. Our chosen strategy achieves the highest SR (19.15%), indicating that compelling the model to articulate a user-facing intent acts as an effective reasoning scaffold. Critically, formats trained without this textual recommendation (Function only and Think + Function) exhibit a catastrophic failure in safety, with False Prediction (FTR) rates near 100%. This demonstrates that generating the intent is indispensable for teaching the model the crucial skill of when not to act.
The study also reveals a crucial trade-off between SR and safety. While adding a Think step (Think + Recommendation + Function) slightly lowered SR, it drastically enhanced safety, slashing the FTR rate to a remarkable 2.21%. This highlights the Think step as a promising direction for building maximally safe agents. Nevertheless, our primary Recommendation + Function approach offers the best-performing balance between SR and reliability, thus validating our core design choice. Further ablation studies, including an analysis of the impact of different contextual dimensions, are detailed in Appendix.
5 Conclusion
In this work, we address the critical bottleneck hindering the transition of mobile agents from a reactive to a proactive paradigm: the lack of a executable, objective, and realistic benchmark. We introduce ProactiveMobile, a comprehensive benchmark that formalizes the proactive task around a four-dimensional context model, incorporates multi-answer annotations, and uniquely mandates an executable function-call sequence output. Our extensive experiments validate that proactivity is a specialized, learnable capability. This is consistently demonstrated as fine-tuning on our benchmark boosts different models’ performance, with our top-performing model achieving a 19.15% success rate—establishing a new state-of-the-art that surpasses even leading proprietary models like o1 (15.71%). This proves the efficacy of ProactiveMobile as an essential tool for targeted training and highlights the significant gap in current models’ out-of-the-box abilities.
While our work establishes a new SOTA, the modest absolute success rates underscore that proactive intelligence is a profoundly challenging research problem, opening up several promising future directions. Key priorities include enhancing models’ multimodal reasoning to close the significant performance gap between text and multimodal tasks, and exploring advanced training methodologies like reinforcement learning for more robust decision-making. Furthermore, our ablation study on output formats reveals a rich trade-off between success rate and safety, warranting deeper investigation into creating agents that are not only effective but also trustworthy. By providing a foundational and challenging testbed, ProactiveMobile aims to catalyze these future innovations, steering the community toward the development of truly intelligent, anticipatory agents.

### v2

@@
 1 Introduction
-Figure 1: A comparison of proactive and reactive paradigm in mobile agents.
+Figure 1: A comparison of proactive and reactive paradigms in mobile agents.
 Fueled by rapid advancements in MLLMs [47, 49, 3], mobile agents have achieved substantial breakthroughs[46, 14] such as interface comprehension [12, 55], conversational interaction [36, 21], and task planning [8, 37].
 However, these agents share a fundamental constraint: they are confined to a reactive paradigm, functioning as passive executors of direct user commands [34, 10].
-These models places the entire cognitive burden on the user, from need identification to goal articulation [30], thereby relegating the agent to the role of a high-level tool and fundamentally limiting its potential for seamless integration into daily life [23].
+These models place the entire cognitive burden on the user, from need identification to goal articulation [30], thereby relegating the agent to the role of a high-level tool and fundamentally limiting its potential for seamless integration into daily life [23].
 The limitations of the reactive paradigm are becoming a critical bottleneck, propelling a fundamental shift towards proactive intelligence—the undisputed next frontier for mobile agents.
@@
 To establish baselines on our benchmark, we fine-tuned Qwen2.5-VL-7B-Instruct [3] and MiMo-VL-7B-SFT-2508 [35] on the training set.
-We then evaluated its performance on ProactiveMobile alongside a suite of leading closed-source models, including o1 [17], GPT-5 [29], and Gemini-2.5-Pro [7].
+We then evaluated their performance on ProactiveMobile alongside a suite of leading closed-source models, including o1 [17], GPT-5 [29], and Gemini-2.5-Pro [7].
 Our fine-tuned Qwen2.5-VL-7B-Instruct achieves a success rate of 19.15% on the exact function sequence matching, significantly outperforming closed‑source models, including o1 (15.71%), GPT-5 (7.39%), and Gemini-2.5-Pro (8.91%).
@@
 LLMs excel at decomposing high-level natural language commands into a series of executable actions.
-However, methods based on static prompting often struggled with long-horizon tasks and dynamic environments [51, 34, 42, 40].
-Some researches explore fine-tuning or reinforcement learning to enhance the reasoning and prediction capabilities of MLLMs in relative tasks [39, 53, 28, 27, 13].
-The maturation of the field is also marked by comprehensive benchmarks [25, 50, 54, 55, 53], which provide standardized environments for evaluating agent performance on realistic tasks.
+However, methods based on static prompting often struggle with long-horizon tasks and dynamic environments [51, 34, 42, 40].
+Some research explores fine-tuning or reinforcement learning to enhance the reasoning and prediction capabilities of MLLMs in related tasks [39, 53, 28, 27, 13].
+The maturation of the field is also marked by comprehensive benchmarks [25, 50, 54, 53], which provide standardized environments for evaluating agent performance on realistic tasks.
 2.2 Proactive Agents
@@
 3.1 Task Definition
-We define proactive intelligence as the task of predicting user’s latent intentions based on their user profile, device status, world information, and behavioral trajectories.
+We define proactive intelligence as the task of predicting users’ latent intentions based on their user profile, device status, world information, and behavioral trajectories.
 The details of these four categories are as follows:
@@
 GUI-Odyssey [25] A dataset for cross-app mobile GUI navigation 21,669
-AITZ [50] Largest dataset in android GUI navigation field 9,413
+AITZ [50] Largest dataset in the Android GUI navigation field 9,413
 CAGUI [53] A real-world Chinese Android GUI benchmark 3,196
 MobileAgentBench Self-collected GUI data from mobile devices 12,481
-Table 1: Summary of GUI dataset.
+Table 1: Summary of GUI datasets.
 3.2.2 Generation Pipeline
@@
 Our chosen strategy achieves the highest SR (19.15%), indicating that compelling the model to articulate a user-facing intent acts as an effective reasoning scaffold.
-Critically, formats trained without this textual recommendation (Function only and Think + Function) exhibit a catastrophic failure in safety, with False Prediction (FTR) rates near 100%.
+Critically, formats trained without this textual recommendation (Function only and Think + Function) exhibit a catastrophic failure in safety, with False Trigger Rate (FTR) rates near 100%.
 This demonstrates that generating the intent is indispensable for teaching the model the crucial skill of when not to act.
 The study also reveals a crucial trade-off between SR and safety.
-While adding a Think step (Think + Recommendation + Function) slightly lowered SR, it drastically enhanced safety, slashing the FTR rate to a remarkable 2.21%.
+While adding a Think step (Think + Recommendation + Function) slightly lowered SR, it drastically enhanced safety, slashing the FTR rate to a 2.21%.
 This highlights the Think step as a promising direction for building maximally safe agents.
@@
 5 Conclusion
-In this work, we address the critical bottleneck hindering the transition of mobile agents from a reactive to a proactive paradigm: the lack of a executable, objective, and realistic benchmark.
+In this work, we address the critical bottleneck hindering the transition of mobile agents from a reactive to a proactive paradigm: the lack of an executable, objective, and realistic benchmark.
 We introduce ProactiveMobile, a comprehensive benchmark that formalizes the proactive task around a four-dimensional context model, incorporates multi-answer annotations, and uniquely mandates an executable function-call sequence output.
@@
 This is consistently demonstrated as fine-tuning on our benchmark boosts different models’ performance, with our top-performing model achieving a 19.15% success rate—establishing a new state-of-the-art that surpasses even leading proprietary models like o1 (15.71%).
-This proves the efficacy of ProactiveMobile as an essential tool for targeted training and highlights the significant gap in current models’ out-of-the-box abilities.
+This demonstrates the efficacy of ProactiveMobile as an essential tool for targeted training and highlights the significant gap in current models’ out-of-the-box abilities.
 While our work establishes a new SOTA, the modest absolute success rates underscore that proactive intelligence is a profoundly challenging research problem, opening up several promising future directions.